# Advancing Image Super-resolution Techniques in Remote Sensing: A Comprehensive Survey

### v2

@@ 7.1 Qualitative Evaluation
  • 1.
    Visual Comparison: Super-resolution images are displayed side by side with original images, HR reference images or other algorithm results to observe the reconstruction quality from the aspects of texture, edge sharpness, structural fidelity, etc.
  • 2.
    Expert visual evaluation: Experts with remote sensing interpretation experience make comprehensive judgments from the perspectives of object recognition, structure preservation, and edge integrity. This method is particularly suitable for the identifiable analysis of objects such as roads, buildings, and farmland boundaries in HR images.
  • 3.
    Application-guided evaluation: Super-resolution results participate in downstream tasks such as land cover classification, object detection, and change detection, etc. Furthermore, the visual quality and semantic preservation are indirectly evaluated through application performance. In addition, the analysis of some typical problems, such as color consistency, artifacts, and over-smoothing may also be involved. This method is highly subjective, and different people may have different standards when checking super-resolution images. To the best of our knowledge, there is not a good solution for these issues currently.
@@ 7.2 Quantitative Evaluation
Quantitative evaluation methods objectively check the quality of reconstructed images through metric values, which provides a unified standard for performance comparison among different algorithms. Many evaluation metrics, such as PSNR and SSIM and information entropy (EN), have been proposed. Specifically, PSNR measures the error between the reconstructed image and the reference image at the pixel level, while SSIM focuses more on the preservation of structural information, which is particularly suitable for the evaluation of highly structural targets such as building outlines and road networks in RSIs. In addition, EN is used to evaluate the richness of spatial information in super-resolution images.
@@
On the other hand, some non-reference metrics, such as NIQE and PI, have also been proposed. They quantify the visual naturalness and quality degradation based on the statistical characteristics of the image, and are particularly suitable for real remote sensing scenes without HR ground truth. In current RSISR research, although quantitative evaluation provides a standardized and repeatable approach, they still face many challenges. Specifically, different evaluation metrics and test data sets are often used in different studies, which leads to a lack of consistency in experimental settings and result interpretation among various methods, making fair comparisons across methods difficult. In addition, each metric can usually only evaluate one aspect of image quality. Therefore, building a comprehensive and unified evaluation framework has become one of the most critical issues in the field of RSISR.
@@ 8 Future Prospects
8.1 Better Evaluation Metrics
Most RSISR methods are evaluated qualitatively by using visual performance and quantitatively by using evaluation metrics in existing papers. Existing metrics usually, such as PSNR and SSIM, can usually only evaluate one aspect of the super-resolved image, and different indicators are used for performance evaluation on different test datasets in different papers, making it difficult to compare algorithms fairly. Therefore, it is desirable to have better metrics. Ideal metrics should be consistent with visual performance and strongly correlated with performance evaluation on downstream tasks. In addition, it is also encouraging to propose evaluation metrics that adapt to diverse remote sensing data sources in the future.
8.2 Better Benchmarks
Although some benchmarks for RSISR have been proposed, the research on RSISR benchmark is still at an early stage. In 2023, Kowaleczko et al. [322] proposed a new benchmark, MuS2, for RSISR. However, the heterogeneity of MuS2 data leads to the inconsistencies between geometric and spectral , ground feature variations caused by large time spans, and inconsistencies between indicators and real perception. These challenges have greatly weakened the ability of the benchmark to fairly evaluate algorithm performance and slowed down the development of SR technology. In addition, Aybar et al. [320] proposed the OpenSR-test benchmark for optical RSISR. Although OpenSR-test focuses on same-day imaging and manual cleaning, the size of its dataset is still limited by the HR image acquisition and fine registration process. A more recent work is the RRSISR benchmark proposed by Wang et al. [323]. Although these benchmarks have made significant progress in real data collection, multi-temporal pairing, and evaluation mechanisms, these datasets rely on heterogeneous sensors, which makes it difficult to unify imaging geometry, spectral response, time span, and band consistency, resulting in insufficient evaluation generalization ability.
Furthermore, current mainstream metrics such as PSNR, SSIM and LPIPS fail to effectively measure multispectral consistency, structural fidelity and practical application performance. In addition, existing benchmarks face challenges in data acquisition costs and scale expansion, and it is difficult to meet the universality and scalability of large-scale, multi-task, and multi-source remote sensing. For this reason, it is urgent to build a benchmark for RSISR with authenticity, universality, and multi-task jointness to promote the model from synthetic optimization to real application.
8.3 Diffusion-Mammba Based RSISR Methods
@@
Diffusion models and Mamba have demonstrated excellent performance in many computer vision tasks. However, the application of such models in RSISR is at a very early stage. The diffusion model generates high-fidelity details through iterative denoising, but its computational efficiency is low. In contrast, the Mamba model has high computational efficiency due to its linear complexity long-range modeling ability, but it still needs to be optimized in recovering high-frequency details. In the future, we expect many RSISR methods based on the diffusion model and Mamba to emerge. Specifically, it is interesting to develop a hybrid architecture of diffusion and Mamba to synergize the advantages of both to improve application performance.
8.4 Application-Oriented RSISR Methods
One of the motivations of RSISR is to improve the performance of practical applications. However, it can be seen from our review that most studies ignore the deep requirements of downstream tasks for image quality. Specifically, existing methods generally use general metrics such as PSNR and SSIM, which may lead to suboptimal performance of downstream tasks. In the future, it is better to focus RSISR research on the application-oriented paradigm. To this end, we expect that the application-oriented RSISR methods will become the mainstream methods in this field. Specifically, it is desirable to conduct degradation modeling with application-oriented, downstream task-embedded evaluation and lightweight adaptive architectures.
8.5 Improving RSISR Efficiency
The efficiency of the RSISR algorithm greatly affects the deployment of real-time applications. Most studies significantly reduce the computational overhead through dynamic computing strategies and lightweight architectures, but still face the trade-off between detail recovery and efficiency. In the future, it is desirable to combine the local perception of CNN with the global modeling capabilities of Mamba or Transformer to reduce computation while maintaining accuracy.

### v3

@@ 7.1 Qualitative Evaluation
  • Visual Comparison: Super-resolution images are displayed side by side with original images, HR reference images, or other algorithm results to observe the reconstruction quality from the aspects of texture, edge sharpness, structural fidelity, etc.
  • Expert visual evaluation: Experts with remote sensing interpretation experience make comprehensive judgments from the perspectives of object recognition, structure preservation, and edge integrity. This method is particularly suitable for the identification analysis of objects such as roads, buildings, and farmland boundaries in HR images.
  • Application-guided evaluation: Super-resolution results participate in downstream tasks such as land cover classification, object detection, and change detection, etc. Furthermore, the visual quality and semantic preservation are indirectly evaluated through application performance. In addition, the analysis of some typical problems, such as color consistency, artifacts, and over-smoothing may also be involved. This method is highly subjective, and different people may have different standards when checking super-resolution images. To the best of our knowledge, there is no good solution for these issues currently.
@@ 7.2 Quantitative Evaluation
Quantitative evaluation methods objectively check the quality of reconstructed images through metric values, which provide a unified standard for performance comparison among different algorithms. Many evaluation metrics, such as PSNR, SSIM, and information entropy (EN), have been proposed. Specifically, PSNR measures the error between the reconstructed image and the reference image at the pixel level, while SSIM focuses more on the preservation of structural information, which is particularly suitable for the evaluation of highly structural targets such as building outlines and road networks in RSIs. In addition, EN is used to evaluate the richness of spatial information in super-resolution images.
@@
On the other hand, some non-reference metrics, such as NIQE and PI, have also been proposed. They quantify the visual naturalness and quality degradation based on the statistical characteristics of the image, and are particularly suitable for real remote sensing scenes without HR ground truth. In current RSISR research, although quantitative evaluation provides a standardized and repeatable approach, it still faces many challenges. Specifically, different evaluation metrics and test data sets are often used in different studies, which leads to a lack of consistency in experimental settings and result interpretation among various methods, making fair comparisons across methods difficult. In addition, each metric can usually only evaluate one aspect of image quality. Therefore, building a comprehensive and unified evaluation framework has become one of the most critical issues in the field of RSISR.
8 Recent Development Characteristics
In 2015, CNN was first applied to SR tasks, and the milestone work SRCNN [32, 33] was proposed. Further, msiSRCNN [151] was proposed in 2016 for RSISR. Since then, various CNN-based RSISR methods [54, 61, 63] have emerged. In addition, GANs [201] and Transformer [232] were introduced into the RSISR field in 2018 and 2021, respectively. These architectural models have been widely studied and used in the field of RSISR. In 2022, diffusion models [252] were also introduced into RSISR, and many diffusion model-based RSISR methods have been proposed. Another new architecture is Mamba, which was introduced into RSISR in 2024. We anticipate that an increasing variety of deep learning models will be adopted for RSISR tasks in order to further enhance performance.
8.2 Most RSISR methods are supervised
Most RSISR methods involving deep learning are supervised, because RSI usually has the characteristics of high geometric alignment accuracy, which means that paired training samples can be easily constructed by downsampling. In addition, mainstream evaluation metrics such as PSNR and SSIM require ground truth (i.e., real HR image) as a reference, which further demonstrates the advantages of supervised learning. Although unsupervised and self-supervised methods have emerged in recent years, they still face great limitations in the field of remote sensing, such as poor sample diversity and a lack of transferable pre-training. Therefore, most RSISR methods still rely on the supervised paradigm to obtain stable reconstruction performance.
8.3 RSISR methods for different types of images
Most existing RSISR methods perform HR reconstruction on optical RSI, which is mainly collected from satellite images such as WorldView, Gaofen, and Sentinel-2. Recently, some researchers started developing methods to conduct SR on different types of RSI. However, it is not uncommon to perform SR on different types of RSI. For example, in 2012, He et al. [20] proposed an SR method for SAR images. Specifically, they achieved SR reconstruction on a very small test set by combining multi-dictionary and sparse coding spatial pyramid machine training. Similar work includes RSMAP [64] and Guo et al. [114].
Recently, many SR methods for HSI and MSI have been proposed. These images contain rich spectral information in a large number of bands and have broad research prospects. For example, from 2021 to 2024, many HSISR methods were proposed, covering Transformer-based [79, 86], Diffusion-based [85, 94], and self-supervised learning methods [40, 75, 90]. This diversity of RSISR tasks reflects the characteristics of different types of images, such as noise in SAR, spectral fidelity in HSI, and temporal consistency in satellite images, which require customized solutions. However, it is desirable to develop a general SR framework for different types of RSI, which may face great challenges in model generalization, cross-modal knowledge transfer, and computational efficiency.
8.4 Research on benchmarks
Although RSISR has made significant progress in recent years, benchmark research for RSISR still faces challenges. Unlike many basic tasks in computer vision, there are few studies on RSISR benchmarks, and there is a lack of general and high-quality RSISR benchmark datasets, which makes model performance comparisons unstable. In addition, most existing benchmarks [386, 387] rely on idealized downsampling degradation models, which limit the real-world adaptability of the methods.
Until recently, some researchers have started to create benchmarks in the field of RSISR. For example, Kowaleczko et al. [388] proposed the first benchmark dataset MuS2 based on real multi-temporal LR images and HR reference images, which breaks through the limitations of traditional simulated data and enhances the practicality of RSISR. In addition, Wang et al. [389] constructed a real degradation benchmark dataset, RRSISR, based on spectral camera imaging to address the mismatch between simulated degradation and real scenes in RSISR, and proposed a reference table-based block exit (RPE) method to improve efficiency through dynamic calculation.
8.5 Application-oriented RSISR methods
Most existing RSISR methods do not consider downstream applications in the process of image reconstruction. Therefore, these RSISR methods generate visually pleasing reconstructed images, but this may be suboptimal for downstream tasks. This leads to a growing gap between metric-based performance (e.g., PSNR, SSIM) and actual utility in downstream applications. Bridging this requires RSISR models aligned with task-specific objectives, not just perceptual quality.
In the past three years, some studies have started deeply coupling the SR process with specific remote sensing application tasks to achieve a paradigm shift from visual quality-oriented to application-oriented. For example, Zhang et al. [212] proposed an RSI reconstruction method based on SRGAN to construct a high-quality landslide training set from LR images. In addition, Kong et al. [223] proposed a dual RSS-GAN method to integrate high-resolution Planet Fusion data to spatially enhance NDVI and NIRv, significantly improving the temporal expression and estimation accuracy of the surface index. Recently, Sun et al. [237] proposed the RSISR framework based on Transformer and self-supervision mechanism to effectively improve the performance of image interpretation. Fan et al. [218] proposed a GAN-based RSISR method, which significantly improved the accuracy of vegetation index inversion and yield prediction by improving the spatial resolution of ginkgo tree RSIs. Compared with most RSISR methods, these methods directly consider the performance of downstream applications during the image reconstruction process, which provides reconstructed images that are more suitable for specific applications.
8.6 Programming frameworks
@@ 9 Future Prospects
Although the RSISR field has made remarkable progress, significant challenges remain, especially in terms of real-world applications and performance evaluation. In response to the evolving needs of remote sensing technologies, this section expands on emerging directions and the underlying research gaps. Key areas of future development include real-world degradation modeling, the use of foundation models, multi-modal super-resolution (SR), and improvements in model efficiency and evaluation metrics. These emerging trends will define the trajectory of RSISR over the next decade.
9.1 Real-World Degradation Modeling in RSISR
Despite growing recognition of its importance, real-world degradation modeling in RSISR remains at an early stage. Traditional RSISR methods predominantly rely on synthetic degradation models, such as bicubic downsampling or noise addition. However, these simplistic approaches fail to capture the full spectrum of complex and compound distortions encountered in real-world remote sensing scenarios. Degradation processes in this domain typically stem from multi-source factors, including atmospheric scattering effects (e.g., haze, turbulence, simulated by atmospheric point spread functions - PSFs), sensor-specific imperfections (e.g., CCD/CMOS noise patterns, lens blur described by modulation transfer functions - MTFs), geometric distortions from platform instability, radiometric inconsistencies, temporal variations across acquisition cycles (e.g., seasonal illumination changes and phenological variations), and preprocessing artifacts. These distortions interact in non-trivial, often sensor-specific ways.
Consequently, most state-of-the-art RSISR models are trained on idealized datasets synthesized using these inadequate techniques, limiting their real-world applicability. A related work [78] in 2022 introduced the Residual Balanced Attention Network (RBAN), representing initial steps towards addressing this gap. Nevertheless, these efforts remain exploratory. A critical need persists to systematically develop comprehensive degradation models that explicitly incorporate: (1) Physical Priors: Integrating domain knowledge of atmospheric PSFs and sensor MTFs to ground the modeling in physics. (2) Multi-Temporal Dynamics: Accounting for seasonal illumination changes and phenological variations that alter scene characteristics over time. (3) Cross-Sensor Adaptability: Creating unified frameworks capable of handling the distinct characteristics and degradation profiles of optical, SAR, and hyperspectral sensors, which are the critical requirements for operational multi-source Earth observation.
Achieving this will likely involve leveraging multi-source satellite data with adversarial training to generate realistic degradation patterns, alongside unsupervised domain adaptation techniques to enhance robustness against sensor heterogeneity. Future research should prioritize both data-driven approaches, such as learning degradation priors from unpaired HR/LR image distributions, and physics-informed strategies, like modeling atmospheric scattering via radiative transfer simulations. By systematically bridging the simulation-to-reality gap through such advanced modeling, RSISR can achieve significantly improved generalization capabilities. This advancement is crucial for critical applications such as disaster response and precision agriculture.
9.2 Foundation Models and Transfer Learning in RSISR
The remarkable success of foundation models pre-trained on large-scale datasets, like natural language processing and computer vision, prompts a natural question: Can RSISR benefit from similar paradigms? The answer appears promising, albeit with significant challenges. Unlike general vision tasks, RSISR must contend with highly heterogeneous data characterized by variations across sensors, spectral ranges, acquisition conditions, and spatio-temporal resolutions. While foundation models pre-trained on massive natural image datasets (e.g., ImageNet) provide strong general features, they often lack sensitivity to the specific spectral and radiometric nuances inherent in remote sensing imagery.
A more effective approach may involve developing domain-adapted foundation models, pre-trained on large-scale satellite imagery corpora—potentially spanning multiple modalities (optical, SAR, hyperspectral). Subsequently, transfer learning becomes a powerful tool: fine-tuning such a domain-specific foundation model on downstream RSISR tasks (e.g., crop monitoring, disaster assessment) with limited supervision could yield robust, low-data solutions, thereby reducing training data requirements and enabling faster deployment across different domains. This approach allows RSISR models to leverage vast amounts of general computer vision knowledge while efficiently adapting to new sensor types and degradation conditions.
Furthermore, integrating explicit remote sensing priors into foundation architectures, such as atmospheric correction models, sensor-specific calibration, and radiation characteristics, may greatly improve both the accuracy and interpretability of the resulting super-resolved images, enhancing their applicability in real-world scenarios. The future lies in achieving high cross-task, cross-domain transferability, where a unified model trained for RSISR could also effectively support related tasks like change detection, land cover classification, and object recognition with minimal additional training.
9.3 Multi-Modal RSISR Methods
Traditional RSISR methods focus primarily on single-modality inputs, such as typically HR optical (RGB or panchromatic) and SAR imagery. However, remote sensing data is inherently multi-modal, with satellite platforms increasingly offering complementary observations including multispectral, hyperspectral, SAR, LiDAR, and thermal infrared data. These modalities are complementary: optical data provide rich spatial and spectral detail under clear skies, SAR enables all-weather, day-night imaging, and hyperspectral data reveal material-level spectral patterns.
Multi-modal SR techniques that combine information from these various sensors offer significant promise for enhancing image resolution and quality, particularly in scenarios where HR reference data for a single modality are unavailable or incomplete. For instance, LR hyperspectral data could be enhanced using HR panchromatic or multispectral guidance, combining fine spatial structure with detailed spectral signatures. Similarly, HR optical images could be fused with LR radar or thermal data to fill in missing details.
This avenue holds great potential for providing a more comprehensive view of the environment, crucial for applications like precision agriculture, infrastructure monitoring, urban planning, and climate science. However, significant challenges remain, such as the alignment and fusion of multi-modal data complicated by sensor misalignments, different geometries, spectral responses, spatial footprints, acquisition times, and temporal asynchrony.
Future research must therefore focus on addressing these technical hurdles, including cross-modal registration, domain alignment, and developing adaptive fusion strategies. Success may lie in techniques like transformer-based attention mechanisms that learn modality-specific contributions, ultimately enabling the full potential of multi-modal SR for unprecedented accuracy in remote sensing applications.
9.4 Better Evaluation Metrics and Benchmarks
Most RSISR methods are evaluated qualitatively by using visual performance and quantitatively by using evaluation metrics in existing literature. Current mainstream metrics, such as PSNR, SSIM, and LPIPS, remain overly simplistic and problematic. While PSNR and SSIM primarily assess pixel-level accuracy and are useful for basic comparisons, they fail to capture critical aspects essential for remote sensing, such as perceptual realism, application utility, structural and contextual fidelity, spectral consistency, and multi-modal alignment. LPIPS introduces a valuable perceptual assessment but still lacks interpretability in domain-specific tasks. Crucially, these metrics usually evaluate only one aspect of the super-resolved image, and the inconsistent use of different indicators across papers and diverse test datasets makes fair algorithm comparison difficult. Therefore, developing more comprehensive, robust, and standardized evaluation metrics is imperative. Ideal metrics should be consistent with visual performance, strongly correlated with downstream task performance, and capable of assessing multiple dimensions of quality, including spatial resolution, spectral fidelity, structural fidelity, and cross-modal consistency involving semantic alignment across modalities like SAR-optical fusion. Task-aware and perception-aware metrics, such as evaluating structural similarity in object space using object detection/classification accuracy, represent a promising future direction.
In parallel, the existing benchmarks for RSISR, though evolving, are still in their early stages and face significant limitations. Recent contributions like MuS2 [388], OpenSR-Test [386], and RRSISR [389] represent valuable progress in incorporating real data, addressing multi-temporal pairing, and refining evaluation mechanisms. However, existing RSISR benchmarks face significant limitations. The inherent heterogeneity of remote sensing data leads to inconsistencies in geometric and spectral properties, ground feature variations across time, and mismatches between quantitative metrics and human perception. For example, MuS2 suffers from diverse sensor characteristics, while OpenSR-Test, despite its focus on same-day imaging and manual cleaning, is constrained by the high cost of acquiring high-resolution references and precise registration. These factors reduce evaluation consistency and generalization, making it difficult to fairly compare algorithms. Additionally, current benchmarks struggle with data acquisition costs and lack scalability, limiting their applicability to large-scale, multi-task, and multi-source remote sensing scenarios.
Consequently, there is an urgent need for next-generation benchmarks that emphasize authenticity, universality, and task relevance. Such benchmarks should reflect the full complexity of operational environments by integrating multi-temporal, multi-spectral, and multi-sensor data, while explicitly modeling real-world challenges such as cloud cover, atmospheric interference, and sensor noise. Specifically, an ideal benchmarking suite should be modular and support: (1) Multi-source and multi-temporal data integration. (2) Task-embedded evaluations (e.g., linking SR quality to crop classification or damage detection).(3) Realistic degradation conditions. Only by assessing models within such diverse, authentic, and task-aligned contexts using robust, multidimensional metrics can we meaningfully evaluate their practical value and push RSISR toward operational maturity.
9.5 Diffusion-Mamba Based RSISR Methods
@@
Diffusion models and Mamba have demonstrated excellent performance in many computer vision tasks. However, the application of such models in RSISR is at a very early stage. The diffusion model generates high-fidelity details through iterative denoising, but its computational efficiency is low. In contrast, the Mamba model has high computational efficiency due to its linear complexity and long-range modeling ability, but it still needs to be optimized in recovering high-frequency details. In the future, we expect many RSISR methods based on the diffusion model and Mamba to emerge. Specifically, it is interesting to develop a hybrid architecture of diffusion and Mamba to synergize the advantages of both to improve application performance.
9.6 Application Scenarios-Oriented RSISR Methods
A primary motivation for RSISR is to enhance the performance of downstream practical applications. However, a critical limitation persists in current research: most studies overlook the deep-seated quality requirements imposed by specific downstream tasks. This is evident in the prevalent reliance on generic image quality metrics like PSNR and SSIM. These metrics often fail to correlate strongly with the ultimate performance goals of applications such as object detection, land cover classification, or change detection, potentially leading to suboptimal task performance. Consequently, future RSISR research must pivot towards an application-oriented paradigm, establishing it as the mainstream approach in the field.
The necessity for this application scenario-oriented research stems from the vast diversity encountered in real-world scenarios:
(1) Data Heterogeneity: Current research is largely driven by benchmark datasets covering a limited range of scenarios. In contrast, practical RSISR applications involve diverse data from various sensors (optical, SAR, multispectral, hyperspectral), acquisition conditions (atmosphere, illumination, season), and temporal resolutions (daily, weekly, monthly).
(2) Scenario-Specific Requirements: Different application scenarios impose fundamentally distinct quality requirements on super-resolved imagery. Understanding these requirements is critical for selecting or designing appropriate RSISR methodologies:
  • Agricultural Monitoring: Super-resolved images must not only achieve high spatial resolution but, critically, preserve spectral fidelity to accurately differentiate crops, weeds, soil conditions, and stress indicators.
    Method Implication: Incorporating spectral constraints (e.g., spectral loss functions, dedicated network branches) is recommended. Evaluation should emphasize spectral fidelity metrics such as the Spectral Angle Mapper (SAM) and downstream classification accuracy. Typical RSISR methods include ViT-ISRGAN [115], JF-CNNSSR [390], NGSTGAN [391], CSSF [392] and SSU-Net [393], etc.
  • Urban Planning: Accurate reconstruction of fine structural details in buildings, roads, vehicles, and infrastructure is a primary objective.
    Methodological Implication: Techniques that enhance spatial detail fidelity and structural integrity, such as those employing perceptual or adversarial losses, or edge-aware mechanisms, are prioritized. Evaluation should incorporate perceptual quality metrics (e.g., LPIPS) and assess object detection and localization performance. Typical RSISR methods include EGSRN [394], EEGAN [59], CEEGAN [214], SeaNet [395] and SMSR [156], etc.
  • Disaster Management: Temporal consistency across image sequences acquired at different times becomes paramount for reliably tracking event progression, such as flood extent or fire spread.
    Method Implication: Temporal SR or fusion techniques, potentially employing recurrent structures, optical flow, or temporal consistency losses, are crucial. Evaluation must assess temporal stability and change detection accuracy. Typical RSISR methods include ESRGAN [396], MCW-ESRGAN [220], GRNN [8], FireSRnet [397] and SSR-GAN [398], etc.
  • Military/Security Surveillance: This domain necessitates exceptionally high spatial detail for accurate target identification and characterization, demanding precise edge definition and fine texture reproduction. Real-time processing capability may also be required.
    Methodological Implication: Prioritize SR methods that enhance spatial sharpness (e.g., GAN-based approaches). If real-time operation is mandated, incorporate efficiency optimizations (e.g., lightweight models, acceleration techniques). Evaluation should emphasize target detection/recognition metrics and image interpretability. Typical RSISR methods include LRSD-ADMM-NET [399], Joint-SRVDNet [400], MADNet [401], YOLOSR-IST [402] and Liu et al. [403], etc.
  • Environmental Monitoring: This application primarily involves tracking land cover, water quality, and ecosystem dynamics over time. It demands high spectral fidelity for accurate material discrimination alongside moderate spatial detail, with critical emphasis on temporal consistency for robust change analysis.
    Methodological Implication: Techniques that maintain spectral-spatial balance while incorporating multi-temporal data and temporal constraints are optimal. Evaluation must integrate spectral metrics (e.g., SAM, RMSE), change detection accuracy (e.g., F1-score), and temporal consistency measures (e.g., temporal correlation index). Typical RSISR methods include Zhang et al. [138], GSSR [404], CSRNet [405] and STHNN [406], etc.
Table 4: Recommended RSISR Method Considerations for Key Application Scenarios
Application Scenario Core Quality Requirements Recommended RSISR Methods Key Evaluation Metrics
Agricultural Monitoring Spectral Fidelity Moderate Spatial Detail Explicit spectral constraints: Spectral loss (SAM, Spectral MSE), physics-informed models, spectral attention. Prioritize spectral accuracy over sharpness. Recommended methods: ViT-ISRGAN [115], JF-CNNSSR [390], NGSTGAN [391], CSSF [392] and SSU-Net [393], etc. Image Quality: SAM, Spectral RMSE. Task: Crop/weed/soil classification accuracy, stress detection precision.
Urban Planning High Spatial Detail Structural Fidelity (edges, lines) Texture Clarity Enhancing spatial detail: Perceptual loss, adversarial training, edge-aware losses, attention mechanisms. Reconstructing sharp man-made structures. Recommended methods: EGSRN [394], EEGAN [59], CEEGAN [214], SeaNet [395] and SMSR [156], etc. Image Quality: LPIPS, FID (optional), edge sharpness metrics. Task: Building footprint accuracy, road network completeness, vehicle detection & localization precision.
Disaster Management Temporal Consistency Moderate Spatial Detail Fast Processing Temporal SR/Fusion: Utilizing recurrent networks, optical flow, 3D convolutions, explicit temporal consistency losses. Lightweight architectures for rapid response. Recommended methods: ESRGAN [396], MCW-ESRGAN [220], GRNN [8], FireSRnet [397] and SSR-GAN [398], etc. Image Quality: Temporal PSNR/SSIM, temporal stability index. Task: Change detection accuracy (IoU, F1-score), flood/fire spread mapping precision, response time.
Military/Security Surveillance Very High Spatial Detail (small targets) Edge Sharpness Texture Fidelity Real-time (potential) High-performance spatial SR: GAN-based methods for sharpness, methods with strong edge/texture enhancement. Lightweight or accelerated models if real-time is needed. Recommended methods: LRSD-ADMM-NET [399], Joint-SRVDNet [400], MADNet [401], YOLOSR-IST [402] and Liu et al. [403], etc. Image Quality: High-resolution target visibility (subjective/interpretability scales), LPIPS. Task: Small target detection rate, target recognition/identification accuracy, inference speed.
Environmental Monitoring Spectral Fidelity Temporal Consistency Moderate Spatial Detail Methods balancing spectral accuracy and temporal stability: Spectral constraints combined with temporal fusion/consistency mechanisms. Robust to seasonal/illumination variations. Recommended methods: Zhang et al. [138], GSSR [404], CSRNet [405] and STHNN [406], etc. Image Quality: SAM, Temporal stability metrics. Task: Land cover classification accuracy (long-term), change detection accuracy (slow changes like deforestation, erosion), thematic map quality.
Crucially, the ultimate evaluation must be rooted in the downstream application’s performance metrics (e.g., detection rate, classification accuracy, change map quality), moving decisively beyond reliance on generic PSNR/SSIM scores. This application-driven evaluation should also guide the design of loss functions during model optimization. This inherent diversity in data types, application scenarios, and core quality requirements dictates that future RSISR methods must be fundamentally tailored: (1) Application-Driven Degradation Modeling & Evaluation: Degradation models should be developed specifically for target application environments and sensor characteristics. Furthermore, evaluation must move beyond generic metrics (PSNR/SSIM) and incorporate task-embedded performance measures directly linked to the downstream application’s objectives, such as object detectability, change detection accuracy, or classification precision. Optimization and loss functions should align with these end-goal metrics. (2) Adaptive & Lightweight Architectures: Model architectures need to be lightweight for potential deployment on edge devices and adaptive to prioritize the specific quality attributes critical for a given application (e.g., spectral fidelity for agriculture, structural integrity for urban analysis, temporal stability for disaster monitoring). This ensures core requirements are met efficiently.
As a preliminary guidance, Table 4 provides an indicative mapping between representative remote sensing application scenarios and corresponding RSISR method characteristics. While not exhaustive, this mapping summarizes emerging consensus from existing literature and offers a conceptual bridge from application-driven demands to method-level decisions.
It is important to emphasize that these recommendations are indicative rather than prescriptive. The actual method choice should consider additional factors such as data availability, model complexity, hardware constraints, and system latency. Ultimately, the path forward for RSISR lies not merely in matching algorithms to applications post hoc, but in embedding downstream task objectives directly into the SR model training so that the super-resolution process becomes an integral component of the application pipeline rather than a preprocessing step.
In summary, the future of RSISR research lies in deeply understanding and responding to the unique demands of specific application scenarios. By advancing application-driven degradation modeling, adopting task-embedded evaluation frameworks, and designing lightweight adaptive architectures, robust RSISR methods can be developed. These methods will genuinely meet the diverse needs of real-world applications, delivering substantial improvements in the performance of these ultimate tasks.
9.7 Computational Resources and Model Efficiency
The efficiency of RSISR algorithms is increasingly critical as remote sensing datasets continue to grow in size, resolution, and complexity. Unlike general computer vision tasks, RSISR must handle large-scale multi-source data with high spatial resolution, multiple spectral bands, and dense temporal sequences. A single scene may span gigabytes of data, and tasks such as multi-modal fusion or long-term monitoring significantly increase computational demands. This poses clear challenges for both training and inference, especially in real-time or edge-based applications like UAV monitoring and disaster response.
To address these challenges, future RSISR methods should prioritize efficient model designs that reduce resource consumption without sacrificing performance. One promising direction involves lightweight architectures that integrate convolutional layers for local feature extraction with attention mechanisms (e.g., transformers or Mamba models) for global context modeling. These hybrid approaches enable better accuracy–efficiency trade-offs, especially when applied to large-scale or low-latency tasks. Another important direction is adaptive computation, where models dynamically adjust their inference complexity based on scene content or application requirements. Techniques like early-exit mechanisms, spatial pruning, or confidence-based sampling can significantly reduce redundant computation while maintaining output quality. For instance, simpler processing may suffice for homogeneous regions, while more complex areas receive deeper inference.
At the system level, distributed computing and cloud-based infrastructures offer scalable solutions for processing high-volume satellite data. Efficient workload partitioning across temporal or spatial segments enables parallelism during both training and deployment. Leveraging heterogeneous hardware platforms, such as GPU clusters or specialized accelerators, can further reduce latency and improve throughput, especially in large-scale or urgent-response scenarios.
Looking ahead, RSISR systems must be designed with end-to-end efficiency in mind, from sensor-level data handling to inference and downstream integration. By combining compact architectures, adaptive computation strategies, and scalable infrastructure, future models can achieve robust performance across diverse operational contexts, balancing precision with real-world constraints.